1.Find the domain of the function 𝑓(π‘₯)=log4(π‘₯2+4π‘₯βˆ’5𝑒π‘₯βˆ’1).

Β 

2. Construct the tangent line to the graph of the function 𝑓(π‘₯)=√π‘₯+2π‘₯ which is perpendicular to the line π‘₯+3π‘¦βˆ’2=0.

Β 

3. Find the maximal intervals of monotonicity of the function 𝑓(π‘₯)=arctg(π‘₯2βˆ’4π‘₯).

Β 

4. Calculate the integral ∫π‘₯2(π‘₯3+4)2 𝑑π‘₯2βˆ’1.

Β 

5. Find the particular solution of the differential equation 𝑦′=βˆ’(2𝑦+1)β‹…tg(π‘₯) which fulfills the initial condition 𝑦(πœ‹4)=12.

Β 

6. Solve the matrix equation 2𝒳+π’œ=3β„¬βˆ’π’œπ’³ if π’œ=(βˆ’4βˆ’332), ℬ=(1βˆ’121).

Β 


1
Expert's answer
2022-05-27T12:52:58-0400

1.


"x^2+4\ud835\udc65\u22125e^x\u22121>0"

"x<-4.252"

Domain: "(-\infin, -4.252)"


2.


"x+3y-2=0"

"y=-\dfrac{1}{3}x+\dfrac{2}{3}"

The slope of the tangent line

"slope=m=-\dfrac{1}{-\dfrac{1}{3}}=3"

"𝑓(𝑥)=\sqrt{x}+2𝑥"

"f'(x)=\dfrac{1}{2\sqrt{x}}+2"

"slope=m=3=\dfrac{1}{2\sqrt{x}}+2"

"\dfrac{1}{2\sqrt{x}}=1"

"x=\dfrac{1}{4}"

"y(\dfrac{1}{4})=\sqrt{\dfrac{1}{4}}+2(\dfrac{1}{4})=1"

The equation of the tangent line in point-slope form is


"y-1=3(x-\dfrac{1}{4})"

The equation of the tangent line in slope-intercept form is


"y=3x+\dfrac{1}{4}"



3.

Domain:"(-\infin, \infin)"


"f'(x)=(\arctan(x^2-4x))'=\dfrac{2x-4}{1+(x^2-4x)^2}"

"f'(x)=0=>\dfrac{2x-4}{1+(x^2-4x)^2}=0"

"x=2"

If "x<2, f'(x)<0, f(x)" decreases.

If "x>2, f'(x)>0, f(x)" increases.

The function "f" is monotone decreasing on "(-\infin, 2)."

The function "f" is monotone increasing on "(2,\infin)."


4.


"\int x^2(x^3+4)^2 dx"

"u=x^3+4, du=3x^2dx"

"\int x^2(x^3+4)^2 dx=\dfrac{1}{3}\int u^2du=\dfrac{u^3}{9}+C"

"=\dfrac{(x^3+4)^3}{9}+C"

5.


"\dfrac{dy}{2y+1}=-\tan (x)dx"

Integrate


"\int \dfrac{dy}{2y+1}=-\int\tan (x)dx"

"\int\tan (x)dx=\int\dfrac{\sin (x)}{\cos(x)}dx=-\ln|\cos(x)|+\dfrac{1}{2}\ln C"

"\dfrac{1}{2}\ln|2y+1|=\ln|\cos(x)|+\dfrac{1}{2}\ln C"

"2y+1=C\cos ^2 (x)"

Given "y(\dfrac{\pi}{4})=12"


"2(12)+1=C\cos ^2 (\dfrac{\pi}{4})"

"C=50"

The particular solution of the differential equation

"2y+1=50\cos ^2 (x)"

6.


"2X+A=3B-AX"


"A=\begin{pmatrix}
 -4 & -3 \\
 3 & 2
\end{pmatrix}, B=\begin{pmatrix}
 1 & -1 \\
 2 & 1
\end{pmatrix}"

Let


"X=\begin{pmatrix}
 x_{11}& x_{12} \\
 x_{21} & x_{22}
\end{pmatrix}"

Then


"2X=\begin{pmatrix}
 2x_{11}& 2x_{12} \\
 2x_{21} & 2x_{22}
\end{pmatrix}"

"AX=\begin{pmatrix}
 -4 & -3 \\
 3 & 2
\end{pmatrix}\begin{pmatrix}
 x_{11}& x_{12} \\
 x_{21} & x_{22}
\end{pmatrix}"

"=\begin{pmatrix}
 -4x_{11}-3x_{21}& -4x_{12}-3x_{22} \\
 3x_{11}+2x_{21} & 3x_{12}+2x_{22}
\end{pmatrix}"


"2X+AX=\begin{pmatrix}
 -2x_{11}-3x_{21}& -2x_{12}-3x_{22} \\
 3x_{11}+4x_{21} & 3x_{12}+4x_{22}
\end{pmatrix}"

"3B-A=\begin{pmatrix}
 7 & 0 \\
 3 & 1
\end{pmatrix}"

"\begin{pmatrix}
 -2x_{11}-3x_{21}& -2x_{12}-3x_{22} \\
 3x_{11}+4x_{21} & 3x_{12}+4x_{22}
\end{pmatrix}=\begin{pmatrix}
 7 & 0 \\
 3 & 1
\end{pmatrix}"


"-2x_{11}-3x_{21}=7""3x_{11}+4x_{21}=3"

"-2x_{12}-3x_{22}=0""3x_{12}+4x_{22}=1"

"x_{21}=27""-2x_{11}-3x_{21}=7"


"-x_{22}=2""3x_{12}+4x_{22}=1"

"x_{11}=37, x_{21}=-27, x_{22}=-2, x_{12}=3"


"X=\begin{pmatrix}
 37& -27 \\
 3 & -2
\end{pmatrix}"